# Entropy-Lens: The Information Signature of Transformer Computations

### v1

@@ 5.1.3 Entropy profiles identify correct task execution
Then, we applied our pipeline to extract and aggregate and the responses’ entropy profiles and classify them against the correctness of the corresponding LLM-generated answer. We train a k-NN classifier for each LLM and validate it using 10-fold cross-validation. We also conducted a t-test to compare our classifier to a dummy model. This dummy model generates predictions randomly, sampled from a distribution that reflects the proportion of correct and incorrect answers produced by the LLM, ensuring robustness against class imbalance. The results reject the null hypothesis with the k-NN achieving an AUC-ROC between 67.23 and 73.61, depending on prompt type and model (Table 2).
@@
Figure 5: Entropy profiles for ViT model families.
5.2 Entropy-Lens for Vision Transformers
@@
Finally, to demonstrate the versatility and robustness of our approach beyond language modeling, we analyze the entropy profiles of ViTs and DeiTs.
@@
Using 20 classes from ImageNet-1K (Russakovsky et al., 2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 5. We observe that all profiles start with high entropy values, which then decrease, mostly in the final layers. This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1.1), suggesting a universal pattern across domains as different as image processing and natural language processing.
@@
We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between $87$M and $307$M parameters.
Table 2: Performance comparison of different models and prompt types on MMLU accuracy and k-NN AUC-ROC score.
Model Prompt Type LLM Accuracy k-NN AUC-ROC
Llama-3.2-3B-Instruct Base 50.89 73.61 ± 1.52
Humble 58.51 69.90 ± 1.06
Instruct 60.62 67.23 ± 1.62
Gemma-2-2B-it Base 56.10 71.88 ± 1.63
Humble 54.71 72.78 ± 1.15
Instruct 56.38 68.36 ± 1.23
6 Conclusions
In this work, we prototyped a novel model-agnostic interpretability framework for large-scale transformer-based architectures grounded in information-theory. In Section 5.1.1, we showed that the entropy profiles of LLM intermediate predictions identify the LLM’s model family. In Section 5.2 we conduct similar experiments on vision transformers, demonstrating the wide applicability of our framework. Additionally, in Section 5.1.2 we showed that the same entropy profiles can be used to identify the ‘task type’ in LLMs, and in Section 5.1.3 we used them to distinguish between correct and wrong LLM generated answers. Importantly, all our experiments were conducted on frozen off-the-shelf large-scale transformers.
Appendix A Evaluation Details
Table 3: Prompt templates used for TinyStories tasks.
@@ Appendix A Evaluation Details
Table 4: Prompt templates used for the MMLU dataset.

### v2

@@ 5.3 Entropy profiles identify correct task execution
Then, we apply our pipeline to extract and aggregate the responses’ entropy profiles and classify them against the correctness of the corresponding LLM-generated answer. We train a k-NN classifier for each LLM and validate it using 10-fold cross-validation. We also conduct a t-test to compare our classifier to a dummy model. This dummy model generates predictions randomly, sampled from a distribution that reflects the proportion of correct and incorrect answers produced by the LLM, ensuring robustness against class imbalance. The results reject the null hypothesis ($\alpha=0.05$) with the k-NN achieving an AUC-ROC between 67.23 and 73.61, depending on prompt type and model (see Table 3).
@@
Figure 6: PCA projection on aggregated entropy profiles extracted from the topic-format dataset.
5.4 Three regimes of Rényi entropy
Figure 7: Cosine similarity matrices between entropy profiles computed on a subset of the topic-format dataset using different values of $\alpha$ in the Rényi entropy. $\sigma$ denotes the standard deviation of the similarity matrix. Note how similarity flattens for very low and very high $\alpha$, while intermediate values yield more informative profiles.
To qualitatively explore how the Rényi entropy affects the structure of entropy profiles, we compute pairwise cosine similarities between profiles generated with different values of $\alpha$. This analysis is performed on a subset of the topic-format dataset, and the resulting similarity matrices are shown in Figure 7.
We observe three distinct regimes as $\alpha$ varies:
  • For very small values of $\alpha$ (e.g., $\alpha<0.2$), the similarity matrices are nearly flat, with profiles being almost identical across examples. This is expected, as Rényi entropy in this regime weights all tokens with non-zero probability almost equally, and usually all tokens have non-null probability.
  • For large values of $\alpha$ (e.g., $\alpha>20$), the similarity matrices also flatten. In this case, the entropy becomes increasingly dominated by the few tokens with highest probabilities. Since these sets of top tokens tend to have similar cardinalities (in the limit equal to 1), the profiles collapse into a narrow set of values, losing expressiveness and becoming more sensitive to local fluctuations.
  • Between these extremes lies an informative regime—approximately $0.5\leq\alpha\leq 20$—where entropy profiles are heterogeneous enough to retain meaningful variation. This is reflected in the standard deviation of the similarity matrices, which peaks in this interval.
This qualitative observation supports our empirical findings: in Section 5.5, we show that format classification accuracy remains high and stable within this informative $\alpha$ range. Notably, Shannon entropy ($\alpha=1$) falls within this interval, providing a strong justification for its use in the main experiments. By choosing $\alpha=1$, we retain discriminative power while avoiding the need to tune additional hyperparameters.
5.5 Entropy profiles identify text format
We test whether the entropy profiles contain signal about the format of the generated text. In this experiment, we prompt the model to produce short texts across different topics while enforcing one of three predefined formats: poem, scientific piece, or chat log. We call these generated texts the topic-format dataset.
Model $\alpha$ k-NN ROC-AUC
Gemma-2-2b-it 0.5 97.3 ± 1.6
1.0 98.7 ± 1.1
5.0 98.4 ± 1.7
Llama-3.2-1b-it 0.5 97.8 ± 1.6
1.0 97.8 ± 2.4
5.0 96.6 ± 2.6
Table 2: ROC-AUC (one-vs-rest) results of two models with three values of $\alpha$ for the Rényi entropy for format classification. Standard deviations are computed over 10-fold cross-validation splits.
We use entropy profiles computed not only with Shannon entropy ($\alpha=1$), but also with Rényi entropy for $\alpha\in\{0.5,5.0\}$. As shown in Table 2, the k-NN classifier achieves high ROC-AUC scores across all values of $\alpha$, indicating that the format is reliably encoded in the entropy profiles. Moreover, the small variability in performance across $\alpha$ values supports our use of Shannon entropy as a principled, parameter-free default.
To better understand the structure of the entropy profiles, we perform Principal Component Analysis (PCA) and visualize the first two components. As shown in Figure 6, the profiles cluster distinctly by format, forming linearly separable groups in the reduced space. This indicates that format-specific computation patterns are not only detectable by a classifier, but visibly reflected in the global shape of the entropy evolution across layers.
Model Prompt LLM Acc. k-NN AUC-ROC
Llama Base 50.89 73.61 ± 1.52
Humble 58.51 69.90 ± 1.06
Instruct 60.62 67.23 ± 1.62
Gemma Base 56.10 71.88 ± 1.63
Humble 54.71 72.78 ± 1.15
Instruct 56.38 68.36 ± 1.23
Table 3: Performance comparison of two models (Llama-3.2-3B-it and Gemma-2-2B-it) for different prompt types on MMLU accuracy and k-NN AUC-ROC score.
6 Conclusions
In this work, we prototyped a novel model-agnostic interpretability framework for large-scale transformer-based architectures grounded in information-theory. In Section 5.1, we showed that the entropy profiles of LLM intermediate predictions identify the LLM’s model family. Additionally, in Sections 5.2 and 5.5 we showed that the same entropy profiles can be used to identify the ‘task type’ and the ‘format type’ in LLMs, whilst in Section 5.3 we used them to distinguish between correct and wrong LLM generated answers. Moreover, we investigated the validity of using Shannon entropy instead of the more general Rényi entropy, and empirically demonstrated its effectiveness. Importantly, all our experiments were conducted on frozen off-the-shelf large-scale transformers.
6.1 Limitations and Future Work
While this work paves the way to further investigations in information theoretic interpretability, it also presents a number of limitations. First, the concepts of ‘task type’ and ‘format type’ don’t have a formal and well established definition.
Moreover, we showed how different models possess different characteristic entropy profiles. We conjecture that these particular shapes are a byproduct of training procedure and architectural designs, but future research could focus on understanding the precise connections. Currently, we do not have a theoretical explanation for the observed phenomena or their underlying causes.
Another interesting line of research could focus on considering more fine-grained measures of information instead of just an aggregated one such as entropy.
With these limitations in mind, our methodology could be used to probe the reasoning capabilities of LLMs, for instance by comparing the entropy profile of a reasoning task vs. a data retrieval task. If these profiles happen to match, it could be taken as an argument against the ability of LLMs to reason. Conversely, if they do not, it may suggest that some form of reasoning is indeed occurring.
Finally, recent literature explored the use of entropy for private inference (PI), where computations are performed on encrypted data without revealing users’ sensitive information (Jha and Reagen 2025). While previous work focused on the entropy of the attention mechanism, future research could use our methodology to develop PI-friendly applications of LLMs.
Appendix A Appendix Structure
The appendix is organized as follows:
  • Appendix B – Additional Experiments: We assess the generality of our approach by applying it to Vision Transformers (ViTs). This analysis demonstrates that entropy profiles can be extracted and interpreted even in non-language domains, without any modification to our framework.
  • Appendix C – Evaluation Details: We provide full details on the datasets and prompt templates used in our experiments. In particular, we specify how prompts were constructed for task classification, output correctness, and format classification tasks. This section also includes hardware information to support reproducibility.
  • Appendix D – Minimal Implementation: We present a minimal code snippet that reproduces the core entropy profile extraction logic in a few lines of code. While our full codebase offers several optimizations and utilities, this section emphasizes transparency and ease of replication by showcasing the conceptual simplicity of our approach.
Appendix B Additional Experiments
To further test the generality and flexibility of our methodology, we conduct additional experiments beyond the core settings presented in the main text. In particular, we explore how Entropy-Lens performs in a different modality: computer vision. Without any architectural adjustment or fine-tuning, we apply the same framework to Vision Transformers (ViTs) and demonstrate that entropy profiles extracted from visual models exhibit meaningful and interpretable structure. This supports the claim that our method is not only scalable and model-agnostic, but also transferable across modalities.
B.1 Entropy-Lens for Vision Transformers
@@
To demonstrate the versatility and robustness of our approach beyond language modeling, we analyze the entropy profiles of ViTs and DeiTs.
@@
Using 20 classes from ImageNet-1K (Russakovsky et al. 2015), with 20 images per class, and without any modifications to our framework, we generate the entropy profiles shown in Figure 8. We observe that all profiles start with high entropy values, which then decrease, mostly in the final layers. This behavior is qualitatively similar to that of GPTs or larger LLaMa models (Section 5.1), suggesting a universal pattern across domains as different as image processing and natural language processing.
@@
We hypothesize a phase transition in entropy behavior as model size increases, occurring somewhere between 8787M and 307307M parameters.
Figure 8: Entropy profiles for ViT model families.
Appendix C Evaluation Details
This section provides further details on the datasets and prompt templates used to evaluate the effectiveness of entropy profiles in the main experiments. In particular, we describe how we constructed the inputs for three key experimental settings: task type classification using the TinyStories dataset, correctness classification using the MMLU benchmark, and format classification using the topic-format dataset.
In all cases, prompt design plays a critical role in ensuring robust comparisons across experimental conditions. To this end, we employed multiple prompt variations. The subsections below report the full set of templates used, as referenced in Sections 5.2, 5.3, and 5.5 of the main paper.
The scripts used to generate these datasets—along with the full codebase to extract entropy profiles and reproduce all experiments—are shared as part of our code release.
Finally, we also provide information about the hardware used to run our experiments to facilitate reproducibility.
C.1 Prompt templates for TinyStories tasks
In Section 5.2, we describe an experimental setup designed to test whether entropy profiles can identify different types of tasks. To this end, we used the TinyStories dataset (Eldan and Li 2023) and constructed prompts combining short stories with specific task instructions. Each task type—generative, syntactic, and semantic—was associated with two distinct natural language formulations, referred to as task prompts. These are listed in Table 4. By varying the task prompt, we ensure that our classification results are not simply driven by surface-level textual artifacts, but instead reflect deeper computational signatures captured by the entropy profiles.
This table complements the prompt templates (base, reversed, and scrambled) described in the main text and provides the full set of instructions used to elicit different model behaviors.
@@ C.1 Prompt templates for TinyStories tasks
Table 4: Prompt templates used for TinyStories tasks.
C.2 Prompt templates for MMLU correctness classification
In Section 5.3, we test whether entropy profiles can distinguish between correct and incorrect answers produced by language models. To construct the dataset for this experiment, we used the Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al. 2021), applying three distinct prompt styles to elicit different answer behaviors from the models. Table 5 reports the full prompt templates used in this experiment. Each template presents the same multiple-choice question in a different instructional format: the Base format presents the question directly; the Instruct format introduces an explicit instruction to select a single correct answer; and the Humble format includes an additional fallback directive encouraging the model to guess randomly if uncertain.
This variation in prompting allows us to control for instruction framing and to evaluate whether entropy profiles can capture response confidence and correctness robustly across different model behaviors. The table shown here complements the description in the main text.
@@ C.2 Prompt templates for MMLU correctness classification
Table 5: Prompt templates used for the MMLU dataset.
C.3 Prompt Construction for the Topic-Format Dataset
To evaluate whether entropy profiles captures stylistic features of generated text, we constructed a custom dataset referred to as the topic-format dataset. In this setting, models are prompted to generate short texts on various topics, each constrained to adopt one of three specific formats: poem, scientific piece, or chat log. The goal is to determine whether these formats induce distinct entropy profiles, independently of the topic content.
We generated prompts by pairing 150 distinct topics with the following three format instructions:
  • Write a poem about ...
  • Write a scientific piece about ...
  • Simulate a chat log about ...
Each prompt is constructed by concatenating a format prefix with a randomly selected topic (e.g., Write a poem about a planet). The resulting dataset contains 450 prompt completions per model, each paired with its entropy profile computed using Rényi entropy for $\alpha\in\{0.5,1.0,5.0\}$.
All generations were performed using a maximum generation length of 256 tokens. We then segmented the output into 8 equal-length windows and computed an entropy profile for each. The resulting data were stored with format labels and used in the classification and visualization tasks discussed in Section 5.5 of the main paper.
This setup enables robust testing of the extent to which entropy profiles encode formatting cues, beyond topical content or task semantics.
C.4 Experimental and hardware setup
All experiments were conducted on a compute node equipped with an NVIDIA L40 GPU, an Intel Xeon Gold CPU, 128 GB of RAM, and running Ubuntu 22.04. The primary software frameworks used were PyTorch, Transformer-Lens, and HuggingFace Transformers. Inference on LLMs was performed using float16 precision for improved efficiency.
Appendix D Minimal Implementation
To maximize reproducibility and transparency, we provide a minimal implementation of our framework. While our full codebase includes optimizations and utility functions to streamline analysis across models and datasets, the core idea behind Entropy-Lens is conceptually simple and can be expressed in just a few lines of code.
This section presents a compact example that computes the entropy profile of generated tokens using an off-the-shelf language model. Despite its brevity, this snippet captures the essence of our method: extracting intermediate representations, mapping them to vocabulary distributions, and computing their entropies.
D.1 Minimal Entropy Profile Extraction
The code in listing 1 demonstrates how to compute an entropy profile for a single prompt using a standard decoder-only transformer. It relies only on model forward passes and the use of logit-lens-style decoding. No gradients, fine-tuning, or access to internal model weights are required.
1from transformers import AutoTokenizer, AutoModelForCausalLM
2import torch
3
4# Load GPT-2 and set up
5tokenizer = AutoTokenizer.from_pretrained(’gpt2’)
6model = AutoModelForCausalLM.from_pretrained(’gpt2’, device_map="auto").eval()
7tokenizer.pad_token = tokenizer.eos_token
8
9# Define entropy computation
10ln, U = model.transformer.ln_f, model.lm_head
11entropy = lambda x: -torch.sum(x * torch.log(x + 1e-15), dim=-1)
12
13# Prepare input
14input_text = ’The concept of entropy’
15inputs = tokenizer.encode(input_text, return_tensors="pt").to(model.device)
16
17# Generate with hidden states
18outputs = model.generate(inputs,
19 do_sample=True,
20 max_new_tokens=32,
21 output_hidden_states=True,
22 return_dict_in_generate=True,
23 pad_token_id=tokenizer.pad_token_id
24)
25
26# Stack hidden activations and compute entropy signature
27activations = torch.vstack([
28 torch.vstack(h).permute(1, 0, 2) for h in outputs.hidden_states
29])
30entropy_signature = entropy(U(ln(activations)).softmax(dim=-1))
Listing 1: A minimal Python implementation of Entropy-Lens using Huggingface models and Pytorch.